### v1

@@ 2.1 Complexity analysis
Lemma 3.
Suppose that Assumption 1 holds. Let $\{z_{k}\}$ and $\{\hat{z}_{k}\}$ be sequences generated by Algorithm 1. Set $\lambda_{k+1}=\frac{c_{k}}{H_{k}\|z_{k+1}-\hat{z}_{k}\|}$ with $c_{k}$ being a constant. If $\frac{1}{33}\leqslant c_{k}\leqslant\frac{1}{13}$, then for any $s\geqslant 1$, we have $H_{s}\leqslant 2\rho$ and
$\sum_{k=1}^{s}\lambda_{k+1}\geqslant\frac{s^{\frac{3}{2}}}{66\sqrt{3}H_{s}\|z_{0}-z^{*}\|},$ (24)
where $z^{*}=(x^{*},y^{*})$ is an optimal solution of (1).
Proof.
By Step 2(c) in Algorithm 1, we have $H_{k}=H_{k,i}$ at the $k$-th iteration. Then by (10) and Assumption 1, we have
$\frac{H_{k,i-1}}{2}\|z_{k,i-1}-\hat{z}_{k}\|^{2}\leqslant\|F(z_{k,i-1})-F(\hat{z}_{k})-DF(\hat{z}_{k})(z_{k,i-1}-\hat{z}_{k})\|\leqslant\frac{\rho}{2}\|z_{k,i-1}-\hat{z}_{k}\|^{2},$
which means $H_{k}=H_{k,i}=2H_{k,i-1}\leqslant 2\rho$. Note that $H_{k}\leqslant H_{k+1}$ by Step 2 in Algorithm 1. By further combining with the choice that $\lambda_{k+1}H_{k}\|z_{k+1}-\hat{z}_{k}\|\geqslant\frac{1}{33}$, we have
$\big{(}\frac{1}{33H_{s}}\big{)}^{2}\sum_{k=1}^{s}(\lambda_{k+1})^{-2}$ $\leqslant\sum_{k=1}^{s}(\lambda_{k+1})^{-2}\big{(}\frac{1}{33H_{k}}\big{)}^{2}\leqslant\sum_{k=1}^{s}\|z_{k+1}-\hat{z}_{k}\|^{2}.$
@@
$s=\sum_{k=1}^{s}1=\sum_{k=1}^{s}\big{(}(\lambda_{k+1})^{-2}\big{)}^{\frac{1}{3}}(\lambda_{k+1})^{\frac{2}{3}}\leqslant\big{(}\sum_{k=1}^{s}(\lambda_{k+1})^{-2}\big{)}^{\frac{1}{3}}\big{(}\sum_{k=1}^{s}\lambda_{k+1}\big{)}^{\frac{2}{3}}.$ (26)
@@
$s\leqslant\big{(}66\sqrt{3}H_{s}\|z_{0}-z^{\star}\|\big{)}^{\frac{2}{3}}\big{(}\sum_{k=1}^{s}\lambda_{k+1}\big{)}^{\frac{2}{3}},$
which completes the proof. ∎
We are now ready to establish the iteration complexity for the LF-CR algorithm. Let $\epsilon>0$ be any given target accuracy, we denote the first iteration index to achieve $\textsc{gap}(\bar{x}_{N},\bar{y}_{N};\beta)\leqslant\epsilon$ by
$N(\epsilon):=\min\{N|~{}\textsc{gap}(\bar{x}_{N},\bar{y}_{N};\beta)\leqslant\epsilon\},$
where $\textsc{gap}(x,y;\beta)$ is defined as in Definition 1 and
$\bar{x}_{N}=\tfrac{1}{\sum_{k=1}^{N}\lambda_{k+1}}\big{(}\sum_{k=1}^{N}\lambda_{k+1}x_{k+1}\big{)},\quad\bar{y}_{N}=\tfrac{1}{\sum_{k=1}^{N}\lambda_{k+1}}\big{(}\sum_{k=1}^{N}\lambda_{k+1}y_{k+1}\big{)}$
with $\beta$ being specified later. We provide an upper bound on $N(\epsilon)$ in the following theorem.
Theorem 1.
Suppose that Assumption 1 holds. Let $\{z_{k}\}$ and $\{\hat{z}_{k}\}$ be sequences generated by Algorithm 1. Set $\beta=7\|z_{0}-z^{*}\|$, $\lambda_{k+1}=\frac{c_{k}}{H_{k}\|z_{k+1}-\hat{z}_{k}\|}$ with $c_{k}$ being a constant. If $\frac{1}{33}\leqslant c_{k}\leqslant\frac{1}{13}$, then
$N(\epsilon)\leqslant\big{(}\frac{7841\sqrt{3}\rho\|z_{0}-z^{*}\|^{3}}{\epsilon}\big{)}^{\frac{2}{3}}+1,$ (27)
where $z^{*}=(x^{*},y^{*})$ is an optimal solution of (1).
Proof.
Under the convex-concave setting, we have
$f(x_{k+1},y)-f(x,y_{k+1})=$ $~{}f(x_{k+1},y)-f(x_{k+1},y_{k+1})+f(x_{k+1},y_{k+1})-f(x,y_{k+1})$
$\leqslant$ $~{}\langle z_{k+1}-z,F(z_{k+1})\rangle.$ (28)
@@
$f(\bar{x}_{N(\epsilon)-1},y)-f(x,\bar{y}_{N(\epsilon)-1})\leqslant$ $~{}\tfrac{1}{\sum_{k=1}^{N(\epsilon)-1}\lambda_{k+1}}\sum_{k=1}^{N(\epsilon)-1}\lambda_{k+1}\big{(}f(x_{k+1},y)-f(x,y_{k+1})\big{)}$
$\leqslant$ $~{}\tfrac{1}{\sum_{k=1}^{N(\epsilon)-1}\lambda_{k+1}}\sum_{k=1}^{N(\epsilon)-1}\lambda_{k+1}\langle z_{k+1}-z,F(z_{k+1})\rangle.$ (29)
By combining (19), (24) and (2.1), we get
$f(\bar{x}_{{N(\epsilon)-1}},y)-f(x,\bar{y}_{{N(\epsilon)-1}})\leqslant\frac{33\sqrt{3}H_{N(\epsilon)-1}\|z_{0}-z\|^{2}\|z_{0}-z^{*}\|}{({N(\epsilon)-1)}^{\frac{3}{2}}}.$ (30)
@@
$\quad\textsc{gap}(\bar{x}_{N(\epsilon)-1},\bar{y}_{N(\epsilon)-1};\beta)$
$=\max_{x\in\mathbb{B}_{\beta}(x^{\star}),y\in\mathbb{B}_{\beta}(y^{\star})}(f(\bar{x}_{N(\epsilon)-1},y)-f(x,\bar{y}_{N(\epsilon)-1}))$
$\leqslant\max_{x\in\mathbb{B}_{\beta}(x^{\star}),y\in\mathbb{B}_{\beta}(y^{\star})}\frac{33\sqrt{3}H_{N(\epsilon)-1}\|z_{0}-z\|^{2}\|z_{0}-z^{*}\|}{(N(\epsilon)-1)^{\frac{3}{2}}}$
$\leqslant\max_{x\in\mathbb{B}_{\beta}(x^{\star}),y\in\mathbb{B}_{\beta}(y^{\star})}\frac{33\sqrt{3}H_{N(\epsilon)-1}(\|z_{0}-z^{*}\|+\sqrt{2}\beta)^{2}\|z_{0}-z^{*}\|}{(N(\epsilon)-1)^{\frac{3}{2}}}$
$=\frac{33\sqrt{3}(7\sqrt{2}+1)^{2}H_{N(\epsilon)-1}\|z_{0}-z^{*}\|^{3}}{(N(\epsilon)-1)^{\frac{3}{2}}}.$
@@
$N(\epsilon)\leqslant\big{(}\frac{33\sqrt{3}(7\sqrt{2}+1)^{2}H_{N(\epsilon)-1}\|z_{0}-z^{*}\|^{3}}{\epsilon}\big{)}^{\frac{2}{3}}+1\leqslant\big{(}\frac{7841\sqrt{3}\rho\|z_{0}-z^{*}\|^{3}}{\epsilon}\big{)}^{\frac{2}{3}}+1$
which completes the proof. ∎
Remark 1.
Note that for any $k\geqslant 1$, we have $H_{0}\leqslant H_{k}\leqslant 2\rho$. Therefore, in the $k$-th iteration, the number of line searches for $H_{k}$ is upper bounded by $\log_{2}\frac{2\rho}{H_{0}}$. Theorem 1 implies that the iteration complexity of the LF-CR algorithm to obtain an $\epsilon$-optimal solution of (1) with respect to the restricted primal-dual gap is upper bounded by $\mathcal{O}(\frac{\rho\|z^{0}-z^{*}\|^{3}}{\epsilon})^{\frac{2}{3}}$.
Compared to the state-of-the-art algorithm in [47], the LF-CR algorithm achieves the same optimal complexity for finding an $\epsilon$-optimal solution to (1) without requiring prior knowledge of the Lipschitz constant, assuming the Hessian is Lipschitz continuous. Although a state-of-the-art algorithm without requiring prior knowledge of the Lipschitz constant was recently proposed in [40] and demonstrated the same optimal complexity, it also requires the assumption of Lipschitz continuous gradients in addition to the Lipschitz continuous Hessians.
@@ 3 A Fully Parameter-Free Cubic Regularization (FF-CR) Algorithm
Although the LF-CR algorithm and the parameter-free algorithm in [40] do not need for prior knowledge of the Lipschitz constants, they still require some other prior knowledge about the problem parameters, such as the predetermined maximum number of iterations $T$, the upper bound of the distance from the initial point $z_{0}$ to the optimal solution $z^{*}$, etc. In this section, we further propose a fully parameter-free cubic regularization (FF-CR) algorithm that does not require any prior knowledge about the problem parameters. We show that the iteration complexity of the proposed algorithm remains optimal in the absence of such information.
Based on the LF-CR algorithm, the proposed FF-CR algorithm further employs an accumulative regularization method to construct the subproblem and uses a “guess-and-check" strategy to estimate $\|z_{0}-z^{*}\|$ step by step, thereby achieving the goal of fully parameter free. Such accumulative regularization method and “guess-and-check" strategy have also been used to design a first-order parameter-free gradient minimization algorithm for convex optimization problem in [16].
Each iteration of the proposed FF-CR algorithm consists of two loops, the outer loop and the inner loop. In the outer loop, the estimate bound $D_{t}$ of $\|z_{0}-z^{*}\|$ iteratively increases. In the inner loop, for a given estimate $D_{t}$, it solves the accumulative regularization subproblem approximately without the prior knowledge of the Lipschitz constants. More detailedly, it performs the following two important algorithmic components at each inner loop:
@@
Step 1: Input $x_{0},y_{0},M_{0},D_{0}$; Set $t=0$.
Step 2: For a given $D_{t}$, solve the accumulative regularized subproblem:
(a): Set $\bar{x}_{0}^{t}=x_{0}^{t}=x_{0},\bar{y}_{0}^{t}=y_{0}^{t}=y_{0}$, $\sigma_{0}^{t}=0$, $H_{0}^{t}=M_{0}^{t}=M_{0}$, $k=1$.
@@
$\min_{x\in\mathbb{R}^{m}}\max_{y\in\mathbb{R}^{n}}f_{k}^{t}(x,y):=f(x,y)+\dfrac{\sigma_{k}^{t}}{2}\|x-\bar{x}_{k}^{t}\|^{2}-\dfrac{\sigma_{k}^{t}}{2}\|y-\bar{y}_{k}^{t}\|^{2}.$ (35)
   Denote the output of the LF-CR algorithm as $(x_{k}^{t},y_{k}^{t},H_{k}^{t})$.
@@
      $M_{k}^{t}=M_{k,i}^{t},\tilde{z}=z_{k,i}^{t}$; otherwise, set $M_{k,i+1}^{t}=2M_{k,i}^{t},i=i+1,$ go to Step
      2(d)(ii).
(e): If $k\geqslant K_{t}$, set $(x_{t+1},y_{t+1},M_{t+1})=(x_{k}^{t},y_{k}^{t},M_{k}^{t})$; otherwise, set $k=k+1$, go
   to Step 2(b).
Step 3: If $\|\nabla f(x_{t+1},y_{t+1})\|\leqslant\epsilon$, stop; otherwise, set $D_{t+1}=4D_{t},t=t+1$, go to Step 2(a).
To the best of our knowledge, the FF-CR algorithm is the first fully parameter-free algorithm for solving minimax optimization problems, without requiring any a priori knowledge about the problem parameters.
In the following subsection, we will establish the iteration complexity of the FF-CR algorithm to obtain an $\epsilon$-optimal solution of (1) with respect to the the gradient norm.

### v2

@@ 2.1 Complexity analysis
Lemma 4
Suppose that Assumptions 1 and 2 hold. Let $\{z_{k,i}\}$, $\{z_{k}\}$ and $\{\hat{z}_{k}\}$ be the sequences generated by Algorithm 1. Assuming that $\|z_{k+1}-\hat{z}_{k}\|\neq 0$, set $\lambda_{k+1}=\frac{c_{k}}{H_{k}\|z_{k+1}-\hat{z}_{k}\|}$ with $c_{k}$ being a constant. If in addition $\frac{1}{33}\leqslant c_{k}\leqslant\frac{1}{13}$, then for any $s\geqslant 1$, we have $H_{s}\leqslant 2\rho$ and
$\sum_{k=1}^{s}\lambda_{k+1}\geqslant\frac{s^{3/2}}{66\sqrt{3}H_{s}\|z_{0}-z^{*}\|},$ (26)
where $z^{*}=(x^{*},y^{*})$ is an optimal solution of (1) and $\rho$ is the local Hessian Lipschitz constant on the compact set $\mathcal{C}:=conv\{z^{*},z_{k,i},z_{k},\hat{z}_{k},i,k=1,2,\cdots\}$, which is the convex hull generated by $z^{*}$ and the iterated point sequence.
Proof  By Assumption 1 and (9), it is easy to see that for any $k,i\geqslant 1$,
$6H_{k,i}\|z_{k,i}-\hat{z}_{k}\|^{3}=$ $-\left\langle z_{k,i}-\hat{z}_{k},DF(\hat{z}_{k})(z_{k,i}-\hat{z}_{k})\right\rangle-\langle F(\hat{z}_{k}),z_{k,i}-\hat{z}_{k}\rangle$
$\leqslant$ $\leavevmode\nobreak\ \|F(\hat{z}_{k})\|\|z_{k,i}-\hat{z}_{k}\|,$
which implies that
$\|z_{k,i}-z^{*}\|\leqslant\|z_{k,i}-\hat{z}_{k}\|+\|\hat{z}_{k}-z^{*}\|\leqslant\sqrt{\frac{\|F(\hat{z}_{k})\|}{6H_{k,i}}}+\|\hat{z}_{k}-z^{*}\|\leqslant\sqrt{\frac{\|F(\hat{z}_{k})\|}{6H_{0}}}+\|\hat{z}_{k}-z^{*}\|.$
Then, by further combining the above relation with (23), it follows that the sequences $\{z_{k,i}\}$, $\{z_{k}\}$ and $\{\hat{z}_{k}\}$ are bounded. By Assumption 2, there exists $\rho>0$ such that for any $(x,y),(x^{\prime},y^{\prime})\in\mathcal{C}:=conv\{z^{*},z_{k,i},z_{k},\hat{z}_{k},i,k=1,2,\cdots\}$, we have
$\|\nabla^{2}f(x,y)-\nabla^{2}f(x^{\prime},y^{\prime})\|\leqslant\rho\|(x-x^{\prime},y-y^{\prime})\|.$ (27)
By induction and Step 2(c) in Algorithm 1, we then show that
$H_{k}\leqslant 2\rho,\forall\leavevmode\nobreak\ k\geqslant 0.$ (28)
Noting that $H_{0}=\frac{\|DF(z_{0})-DF(\tilde{z}_{0})\|}{\|z_{0}-\tilde{z}_{0}\|}$ $(z_{0}\neq\tilde{z}_{0})$, we can easily see that $H_{0}\leqslant\rho$. Suppose (28) holds for k1k-1. Then, at the $k$-th iteration, for the case $i_{k}=1$, we have $H_{k}=H_{k,1}=H_{k-1}\leqslant 2\rho$. For the case $i_{k}>1$, by Step 2(c) and (27), we have
$\frac{H_{k,i_{k}-1}}{2}\|z_{k,i_{k}-1}-\hat{z}_{k}\|^{2}\leqslant\|F(z_{k,i_{k}-1})-F(\hat{z}_{k})-DF(\hat{z}_{k})(z_{k,i_{k}-1}-\hat{z}_{k})\|\leqslant\frac{\rho}{2}\|z_{k,i_{k}-1}-\hat{z}_{k}\|^{2},$
which means $H_{k}=H_{k,i_{k}}=2H_{k,i_{k}-1}\leqslant 2\rho$. Thus, we have $H_{k}\leqslant 2\rho$ and it finishes the induction.
Note that $H_{k}\leqslant H_{k+1}$ by Step 2 in Algorithm 1. By further combining with the choice that $\lambda_{k+1}H_{k}\|z_{k+1}-\hat{z}_{k}\|\geqslant\frac{1}{33}$, we have
$\bigg{(}\frac{1}{33H_{s}}\bigg{)}^{2}\sum_{k=1}^{s}(\lambda_{k+1})^{-2}$ $\leqslant\sum_{k=1}^{s}(\lambda_{k+1})^{-2}\bigg{(}\frac{1}{33H_{k}}\bigg{)}^{2}\leqslant\sum_{k=1}^{s}\|z_{k+1}-\hat{z}_{k}\|^{2}.$
@@
$s=\sum_{k=1}^{s}1=\sum_{k=1}^{s}\big{(}(\lambda_{k+1})^{-2}\big{)}^{1/3}(\lambda_{k+1})^{2/3}\leqslant\bigg{(}\sum_{k=1}^{s}(\lambda_{k+1})^{-2}\bigg{)}^{1/3}\bigg{(}\sum_{k=1}^{s}\lambda_{k+1}\bigg{)}^{2/3}.$ (30)
@@
$s\leqslant\big{(}66\sqrt{3}H_{s}\|z_{0}-z^{\star}\|\big{)}^{2/3}\left(\sum_{k=1}^{s}\lambda_{k+1}\right)^{2/3},$
which completes the proof.
We are now ready to establish the iteration complexity for the LF-CR algorithm to obtain an $\epsilon$-optimal solution of (1) with respect to the restricted primal-dual gap and gradient norm. Let $\epsilon>0$ be any given target accuracy, we denote the first iteration index to achieve $\textsc{gap}(\bar{x}_{N},\bar{y}_{N};\beta)\leqslant\epsilon$ or $\|F(\hat{z}_{N})\|=0$ by
$N(\epsilon):=\min\{N|\leavevmode\nobreak\ \textsc{gap}(\bar{x}_{N},\bar{y}_{N};\beta)\leqslant\epsilon\leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ \|F(\hat{z}_{N})\|=0\},$
where $\textsc{gap}(x,y;\beta)$ is defined as in Definition 2, and
$\bar{x}_{N}=\tfrac{1}{\sum_{k=1}^{N}\lambda_{k+1}}\bigg{(}\sum_{k=1}^{N}\lambda_{k+1}x_{k+1}\bigg{)},\quad\bar{y}_{N}=\tfrac{1}{\sum_{k=1}^{N}\lambda_{k+1}}\bigg{(}\sum_{k=1}^{N}\lambda_{k+1}y_{k+1}\bigg{)}$
with $\beta$ being specified later. In addition, we denote the first iteration index to achieve $\|F(z_{k+1})\|\leqslant\epsilon$ by
[…]
In the following theorem, we provide an upper bound on $N(\epsilon)$ and $\bar{N}(\epsilon)$.
Theorem 1
Suppose that Assumptions 1 and 2 hold. Let $\{z_{k}\}$ and $\{\hat{z}_{k}\}$ be the sequences generated by Algorithm 1. Set $\beta=7\|z_{0}-z^{*}\|$, $\lambda_{k+1}=\frac{c_{k}}{H_{k}\|z_{k+1}-\hat{z}_{k}\|}$ with $c_{k}$ being a constant. If $\frac{1}{33}\leqslant c_{k}\leqslant\frac{1}{13}$, then
$N(\epsilon)\leqslant\left(\frac{7841\sqrt{3}\rho\|z_{0}-z^{*}\|^{3}}{\epsilon}\right)^{2/3}+1,\leavevmode\nobreak\ \bar{N}(\epsilon)\leqslant\frac{156\rho\|z_{0}-z^{*}\|^{2}}{\epsilon}+1,$ (31)
where $z^{*}=(x^{*},y^{*})$ is an optimal solution of (1) and $\rho$ is defined in Lemma 4.
Proof  Under the convex-concave setting, we have
$f(x_{k+1},y)-f(x,y_{k+1})=$ $\leavevmode\nobreak\ f(x_{k+1},y)-f(x_{k+1},y_{k+1})+f(x_{k+1},y_{k+1})-f(x,y_{k+1})$
$\leqslant$ $\leavevmode\nobreak\ \langle z_{k+1}-z,F(z_{k+1})\rangle.$ (32)
@@
$f(\bar{x}_{N(\epsilon)-1},y)-f(x,\bar{y}_{N(\epsilon)-1})\leqslant$ $\leavevmode\nobreak\ \tfrac{1}{\sum_{k=1}^{N(\epsilon)-1}\lambda_{k+1}}\sum_{k=1}^{N(\epsilon)-1}\lambda_{k+1}\big{(}f(x_{k+1},y)-f(x,y_{k+1})\big{)}$
$\leqslant$ $\leavevmode\nobreak\ \tfrac{1}{\sum_{k=1}^{\bar{N}(\epsilon)-1}\lambda_{k+1}}\sum_{k=1}^{N(\epsilon)-1}\lambda_{k+1}\langle z_{k+1}-z,F(z_{k+1})\rangle.$ (33)
According to the definition of $N(\epsilon)$ and Lemma 1, for all $k\leqslant N(\epsilon)-1$, we have $\|z_{k+1}-\hat{z}_{k}\|\neq 0$. By further combining (21), (26) and (2.1), we get
$f(\bar{x}_{{N(\epsilon)-1}},y)-f(x,\bar{y}_{{N(\epsilon)-1}})\leqslant\frac{33\sqrt{3}H_{N(\epsilon)-1}\|z_{0}-z\|^{2}\|z_{0}-z^{*}\|}{({N(\epsilon)-1)}^{3/2}}.$ (34)
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \quad\textsc{gap}(\bar{x}_{N(\epsilon)-1},\bar{y}_{N(\epsilon)-1};\beta)$
$\leavevmode\nobreak\ =\max_{x\in\mathbb{B}_{\beta}(x^{\star}),y\in\mathbb{B}_{\beta}(y^{\star})}(f(\bar{x}_{N(\epsilon)-1},y)-f(x,\bar{y}_{N(\epsilon)-1}))$
$\leavevmode\nobreak\ \leqslant\max_{x\in\mathbb{B}_{\beta}(x^{\star}),y\in\mathbb{B}_{\beta}(y^{\star})}\frac{33\sqrt{3}H_{N(\epsilon)-1}\|z_{0}-z\|^{2}\|z_{0}-z^{*}\|}{(N(\epsilon)-1)^{3/2}}$
$\leavevmode\nobreak\ \leqslant\max_{x\in\mathbb{B}_{\beta}(x^{\star}),y\in\mathbb{B}_{\beta}(y^{\star})}\frac{33\sqrt{3}H_{N(\epsilon)-1}(\|z_{0}-z^{*}\|+\sqrt{2}\beta)^{2}\|z_{0}-z^{*}\|}{(N(\epsilon)-1)^{3/2}}$
$\leavevmode\nobreak\ =\frac{33\sqrt{3}(7\sqrt{2}+1)^{2}H_{N(\epsilon)-1}\|z_{0}-z^{*}\|^{3}}{(N(\epsilon)-1)^{3/2}}.$
@@
$N(\epsilon)\leqslant\left(\frac{33\sqrt{3}(7\sqrt{2}+1)^{2}H_{N(\epsilon)-1}\|z_{0}-z^{*}\|^{3}}{\epsilon}\right)^{2/3}+1\leqslant\left(\frac{7841\sqrt{3}\rho\|z_{0}-z^{*}\|^{3}}{\epsilon}\right)^{2/3}+1,$
which shows the first inequality of (31).
Next, we establish the iteration complexity of Algorithm 1 with respect to the gradient norm. By the definition of $\bar{N}(\epsilon)$ and Lemma 1, for all $k\leqslant\bar{N}(\epsilon)-1$, we have $\|z_{k+1}-\hat{z}_{k}\|\neq 0$. According to (13) and (14), it follows that
$\|F(z_{k+1})\|\leqslant$ $\|F(z_{k+1})-F(\hat{z}_{k})-DF(\hat{z}_{k})(z_{k+1}-\hat{z}_{k})\|+\|F(\hat{z}_{k})+DF(\hat{z}_{k})(z_{k+1}-\hat{z}_{k})\|$
$\leqslant$ $\frac{H_{k}}{2}\|z_{k+1}-\hat{z}_{k}\|^{2}+6H_{k}\|z_{k+1}-\hat{z}_{k}\|^{2}=\frac{13H_{k}}{2}\|z_{k+1}-\hat{z}_{k}\|^{2}.$
Combining the above relation with (22), we have
$\sum_{k=1}^{\bar{N}(\epsilon)-1}\frac{2}{13H_{k}}\|F(z_{k+1})\|\leqslant\sum_{k=1}^{\bar{N}(\epsilon)-1}\|z_{k+1}-\hat{z}_{k}\|^{2}\leqslant 12\|z_{0}-z^{*}\|^{2}.$
Note that $H_{k}\leqslant 2\rho$ for all $k\geqslant 1$. Then by the definition of $\bar{N}(\epsilon)$, we immediately have
$\bar{N}(\epsilon)\leqslant\frac{156\rho\|z_{0}-z^{*}\|^{2}}{\epsilon}+1,$
which completes the proof.
Remark 1
Note that for any $k\geqslant 1$, we have $H_{0}\leqslant H_{k}\leqslant H_{k+1}\leqslant 2\rho$. At the $k$-th iteration, the number of line searches for $H_{k}$ is upper bounded by $i_{k}=1+\log_{2}\frac{H_{k}}{H_{k-1}}$. Then, we have $\sum_{k=1}^{N(\epsilon)}i_{k}=N(\epsilon)+\log_{2}\frac{H_{N(\epsilon)+1}}{H_{0}}\leqslant N(\epsilon)+\log_{2}\frac{2\rho}{H_{0}}.$ Theorem 1 implies that the total iteration complexity of the LF-CR algorithm to obtain an $\epsilon$-optimal solution to (1) with respect to the restricted primal-dual gap and gradient norm is upper bounded by $\mathcal{O}(\rho^{2/3}\|z_{0}-z^{*}\|^{2}\epsilon^{-2/3})$ and $\mathcal{O}(\rho\|z_{0}-z^{*}\|^{2}\epsilon^{-1})$, respectively.
Compared to the Newton-MinMax algorithm in (Lin et al., 2022), the LF-CR algorithm achieves the same complexity for finding an $\epsilon$-optimal solution to (1) without requiring prior knowledge of the Lipschitz constant, assuming that the Hessian is locally Lipschitz continuous. Moreover, the LF-CR algorithm has better iteration complexity with respect to the Lipschitz constant $\rho$ compared to the ASOM method in (Jiang et al., 2024), which has an iteration complexity of $\mathcal{O}\left((\ell^{4/3}\rho^{2/3}\|z_{0}-z^{*}\|^{2/3}+\rho^{2}\|z_{0}-z^{*}\|^{2})\epsilon^{-2/3}\right)$ (resp. $\mathcal{O}\left((\ell\rho\|z_{0}-z^{*}\|+\rho^{2}\|z_{0}-z^{*}\|^{2})\epsilon^{-1}\right)$) in terms of the $\epsilon$-optimal solution of the restricted primal-dual gap (resp. gradient norm).
@@ 3 A Fully Parameter-Free Cubic Regularization (FF-CR) Algorithm
In this section, we further propose a fully parameter-free cubic regularization (FF-CR) algorithm that does not require any prior knowledge about the problem parameters. We show that the proposed algorithm achieves the complexity bound $\mathcal{O}(\epsilon^{-2/3})$ with respect to the gradient norm in the absence of such information.
Based on the LF-CR algorithm, the proposed FF-CR algorithm further employs an accumulative regularization method to construct the subproblem and uses a “guess-and-check” strategy to estimate $\|z_{0}-z^{*}\|$ step by step. Such accumulative regularization method and “guess-and-check” strategy have also been used to design a first-order parameter-free gradient minimization algorithm for convex optimization problem in (Lan et al., 2023).
Each iteration of the proposed FF-CR algorithm consists of two loops, the outer loop and the inner loop. In the outer loop, the estimated bound $D_{t}$ of $\|z_{0}-z^{*}\|$ iteratively increases. In the inner loop, for a given estimate $D_{t}$, it solves the accumulative regularization subproblem approximately without prior knowledge of the Lipschitz constants. More specifically, it performs the following two important algorithmic components at each inner loop:
@@
Step 1: Input $z_{0},\tilde{z}_{0}$; Set $t=0$.
Step 2: Initialize $M_{0}$, $D_{0}$.
  (a): Set $i=1$, $M_{0,i}=\frac{\|DF(z_{0})-DF(\tilde{z}_{0})\|}{\|z_{0}-\tilde{z}_{0}\|}(z_{0}\neq\tilde{z}_{0})$.
  (b): Update $z_{0,i}$ such that it is a solution of the following nonlinear equation:
$F(z_{0})+DF(z_{0})(z-z_{0})+6M_{0,i}\|z-z_{0}\|(z-z_{0})=0.$
  (c): If
$\|F(z_{0,i})-F(z_{0})-DF(z_{0})(z_{0,i}-z_{0})\|\leqslant\frac{M_{0,i}}{2}\|z_{0,i}-z_{0}\|^{2},$
    $M_{0}=M_{0,i}$, $D_{0}=\|z_{0,i}-z_{0}\|$; otherwise, set $M_{0,i+1}=2M_{0,i}$, $i=i+1,$ go to     Step 2(b).
Step 3: For a given $D_{t}$, solve the accumulative regularized subproblem:
  (a): Set $\bar{z}_{0}^{t}=z_{0}^{t}=z_{0}$, $\sigma_{0}^{t}=0$, $H_{0}^{t}=M_{0}^{t}=M_{0}$, $k=1$.
@@
   with $z_{k-1}^{t}$ as the initial point and $H_{k-1}^{t}$ as the initial guess of Hessian Lipschitz    constant. Denote the output of the LF-CR algorithm as $(z_{k}^{t},H_{k}^{t})$.
@@
    $M_{k}^{t}=M_{k,i}^{t},\tilde{z}=z_{k,i}^{t}$; otherwise, set $M_{k,i+1}^{t}=2M_{k,i}^{t},i=i+1,$ go to Step 3(d)(ii).
  (e): If $k\geqslant K_{t}$, set $(z_{t+1},M_{t+1})=(z_{k}^{t},M_{k}^{t})$; otherwise, set $k=k+1$, go to Step 3(b).
Step 4: If $\|F(z_{t+1})\|\leqslant\epsilon$, stop; otherwise, set $D_{t+1}=4D_{t},t=t+1$, go to Step 3(a).
In the following subsection, we will establish the iteration complexity of the FF-CR algorithm to obtain an $\epsilon$-optimal solution of (1) with respect to the gradient norm.